# High-dimensional dynamics in low-dimensional networks

### v4

@@ Results
This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$. This matrix is also sometimes referred to as the “overlap matrix” [11]. Our main result from Appendix B can be stated as follows:
@@
Here, “low-dimensional dynamics” is interpreted to mean that the covariance matrix of ${\bf{z}}(t)$ has an $\mathcal{O}(1)$ number of dominant eigenvalues, equivalently, ${\bf{z}}(t)$ has a small number of dominant principal component directions. Note that this result is equivalent to the following:
@@
Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives.
@@
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}.$ (9)
We take $W_{0}$ to have rank $r=2$ so $W$ has two dominant singular values (Figure 3a) and we fix
@@
High-dimensional dynamics do not require that $W_{0}$ is normal. Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [34]. All normal matrices are EP, but an EP matrix is not necessarily normal. If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied.
@@
$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
@@
The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks with high-dimensional input. As indicated above by Claim 1, the eigenvalue spectrum is also not useful for determining whether strongly low-rank networks produce low-rank suppression.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 12 in Appendix D; compare to Figure 2). However, many networks arising in nature do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
Biased weights and modular networks
@@
If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large. The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix D; see Appendix B in [35] for related results).
Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network. a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure. c) The alignment matrix shows that the network is EP, but not normal. d) An input that is constant within each block (red) is aligned to the low-rank part, but a random input (gray) is not. e) Response magnitude is suppressed for the aligned input relative to the random input. f) The excitatory (positive; red) component of an input balances with the inhibitory (negative; blue) component to produce a much smaller total (gray) component, a widely observed phenomenon in neural circuits.
@@
where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
@@
where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$. Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e). In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24]. As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Figure 16 in Appendix D), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
@@
Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$. As a result, the cancellation mechanism illustrated in Figure 1h manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 4f), a phenomenon that is widely observed in neural recordings [44, 45, 46, 47, 48, 49, 50] and widely studied in computational models [21, 40, 24, 3]. Hence, the widely studied theory of balanced network models can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to a particular modular network structure.
Amplification of disordered perturbations in networks with spatial structure
@@
Figure 5: Low-rank suppression in a network with spatial structure. a) Connectivity matrix, $W$. Connection strength is a Gaussian function of distance. b) Singular values of $W$ demonstrate effective low-rank structure. c) A spatially smooth perturbation (red) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (gray) is not. d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks).
@@
Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 41]. As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a). Connectivity is also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix D). Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix D).
@@
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes. Hence, perhaps surprisingly, strongly low-rank networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
Low-rank suppression and high-dimensional dynamics in an epidemiological network.
@@
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (11)
@@
where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication. Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix D). In [1], it was shown that the dynamics generated by Eq. (11) on this network are effectively low-dimensional.
@@
Figure 6: Low-rank suppression and high-dimensional dynamics in a real epidemiological network. a) The response to a perturbation aligned with the dominant low-rank part is weaker than the response to a perturbation of the same magnitude in a random direction. b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics.
@@
However, Eq. (11) is completely self-contained without any perturbations. Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the school) and through the natural stochasticity of infection.
Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network produces low-rank suppression and high-dimensional responses to high-dimensional perturbations. To test this prediction, we modified the model by adding an external forcing term,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
@@
Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Figure 19b in Appendix D) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.
Discussion
@@
We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices. We also derived conditions under which low-rank networks generate high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
@@
Our results also rely on a linearization of the dynamics around a stable fixed point. The dimensionality of nonlinear network dynamics away from a stable equilibrium is also an important question. However, nonlinear networks can produce any dynamics [29, 30] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks. Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then using bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations. Combining our approaches by incorporating external inputs into their analytical approach is a promising direction for future work.
It is tempting to ascribe low-rank suppression to the presence of an eigenvalue with large, negative real part (as in Figure 1 and Figure 3b–d). A large, negative eigenvalue is well-known to produce a suppressive dynamic. However, the results in Figure 2, Figure 12 in Appendix D, and Claim 1 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression. Low-rank suppression occurs in strongly low-rank networks regardless of their eigenvalue spectrum.
@@
Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equations – except for Figure 6 and Figure 19 in Appendix D – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 6 and Figure 19 in Appendix D were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at

### v5

@@ IV.5 Conditions for low-dimensional dynamics under Assumptions 1–3
This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$. This matrix is also sometimes referred to as the “overlap matrix” [12]. Our main result from Appendix D can be stated as follows:
@@
Note that this result is equivalent to the following:
@@
Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives. We were unable to prove the converse of these claims (that dynamics are low-dimensional whenever $P$ has a small singular value), even though it is consistent with all of our simulations. Proving the converse or finding a counterexample could be a fruitful direction for future research.
Our proof of Claim 1 relied on an assumption that ${\bf{x}}(t)$ varies more slowly than the timescale, $\tau$, of network interactions. In Appendix E, we consider the case where ${\bf{x}}(t)$ is not necessarily slow. In that case, the dimensionality can be broken down into the contribution from different frequency modes. When ${\bf{x}}(t)$ varies at both high and low frequencies, the results in Claim 1 are still valid (see Figure 14 for an example).
@@
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}$ (7)
with $\tau=1$. In these simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d. Gaussian process with correlation time $\tau_{x}=10$. We take $W_{0}$ to have rank $r=2$ so $W$ has two dominant singular values (Figure 3a) and we fix
@@
High-dimensional dynamics do not require that $W_{0}$ is normal. Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called EP matrices [45]. All normal matrices are EP, but an EP matrix is not necessarily normal. If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 1 is satisfied.
@@
$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$
@@
The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks near a stable equilibrium with high-dimensional input. Instead, as Claim 1 shows, the singular values of the recurrent alignment matrix, $P$, are key to determining the dimensionality of dynamics in these networks.
When low-dimensional dynamics are present, they arise from an effectively feedforward mechanism that takes variability from ${\bf{x}}(t)$ and $W_{1}{\bf{z}}(t)$ aligned with the column space of $V$ and maps it to strong variability along the column space of $U$ (see Appendix D for more details). This mechanism is a form of non-normal amplification [26, 27, 28]. Our results add some new results to the literature on non-normal amplification under the conditions in Assumptions 1–3. First, our results show that this form of non-normal amplification requires that $W$ is highly non-EP, not just highly non-normal. See, for example, the example in Figure 3e–g, in which $W_{0}$ is highly non-normal, but EP so that amplification does not occur. Secondly, our results on low-rank suppression show that the amplified responses in the direction of $U$ associated with non-normal amplification are accompanied by suppressed responses approximately aligned to $V$ (see Figure 2e,f). Finally, Claim 1 shows that non-normal amplification is the only way to generate low-dimensional dynamics under Assumptions 1–3.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 9 in Appendix F; compare to Figure 2). However, many networks arising in nature do not have purely random structure. Next, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
V Biased weights and modular networks
Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network. a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure. c) The alignment matrix shows that the network is EP, but not normal. d) An input that is constant within each block (purple) is aligned to the low-rank part, but a random input (green) is not. e) The norm of $\bf{z}$} is more than 13 times smaller in response to the aligned versus random perturbation. f) The excitatory (positive; red) component of mean input (local and external combined) balances with the inhibitory (negative; blue) component to produce a much smaller total input (gray), a widely observed phenomenon in neural circuits. Network dynamics obey Eq. (7).
@@ V Biased weights and modular networks
If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large. The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix F; see Appendix B in [46] for related results).
@@
where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and standard deviation,
@@
where ${\bf{1}}_{e}$ and ${\bf{1}}_{i}$ are indicator vectors for the excitatory and inhibitory populations. Therefore, external inputs, $\bf{x}$}, that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e). In other words, external inputs that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to inputs that are inhomogeneous within one or both populations, consistent with previous work on balanced network models [24, 31, 30, 32]. As predicted, simulations also show high-dimensional responses to high-dimensional stimuli (Figure 16 in Appendix F), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
@@
Averaging over the excitatory and inhibitory populations in this network represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$. As a result, the cancellation mechanism illustrated in Figure 1h manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 4f), a phenomenon that is widely observed in neural recordings [51, 52, 53, 54, 55, 56, 57] and widely studied in computational models [29, 39, 32, 4]. Hence, the widely studied theory of balanced networks can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to the particular modular network structure arising from Dale’s law.
@@
Figure 5: Low-rank suppression in a network with spatial structure. a) Connectivity matrix, $W$. Connection strength is a Gaussian function of distance. b) Singular values of $W$ demonstrate effective low-rank structure. c) A spatially smooth perturbation (purple) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (green) is not. d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks). The norm of $\bf{z}$} is more than 7 times smaller in response to the smooth versus disordered perturbation. Network dynamics obeyed Eq. (7).
VI Amplification of disordered perturbations in networks with spatial structure
@@
Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 40]. As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection weights in $W_{0}$ decay like a Guassian function of the distance, $d\theta$, between nodes computed with periodic boundaries. We assume strong coupling in the sense that individual weights scale like $1/\sqrt{N}$. In this case, only nearby nodes are strongly connected (Figure 5a). Note that stability requires that connections in $W_{0}$ are negative since coupling is strong. Connectivity is also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix F). Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix F).
@@
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes. Hence, perhaps surprisingly, strongly connected networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
VII Low-rank suppression and high-dimensional dynamics in an epidemiological network.
@@
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ Q{\bf{z}}$ (8)
@@
where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication. Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $Q$ is the proportional to the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [58]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix F). In [1], it was shown that dynamics generated by Eq. (8) on this network are effectively low-dimensional.
@@
Figure 6: Low-rank suppression and high-dimensional dynamics in a real epidemiological network. a) The response to a perturbation aligned with the dominant low-rank part is weaker than the response to a perturbation of the same magnitude in a random direction. The distance of $\bf{z}$} from its fixed point is more than 10 times smaller in response to the aligned versus random perturbation. b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics. Network dynamics obeyed Eq. (7).
@@
However, Eq. (8) is completely self-contained without any perturbations. Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network and through the natural stochasticity of infection. Specifically, ${\bf{z}}(t)$ represents a probability of infection. Actual infection events will be random, which can be approximated by adding a stochastic forcing term to Eq. (8). Moreover, students interact with siblings and parents at home (who are not included in the network represented by $Q$), which produces additional external forcing. As a simple model of these perturbations, we modified the model by adding an external forcing term to obtain
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[Q{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$ (9)
Eq. (9) admits a fixed point when ${\bf{x}}(t)=0$. Following the linearization discussed in Section IV.1 and Appendix B, the effective connectivity, $W$, is defined by the Jacobian matrix at the fixed point. We find that, like $Q$, the effective connectivity matrix, $W$, has a small number of dominant singular values, so our theory predicts that this network exhibits low-rank suppression. Moreover, the recurrent alignment matrix, $P$, defined from $W$ satisfies the conditions for high-dimensional dynamics from Claim 1 (Figure 19a in Appendix F). Simulations of Eq. (9) confirm that the model exhibits low-rank suppression (Figure 6a and Figure 19b–d in Appendix F) and high-dimensional dynamics (Figure 6b) in contrast to the model without external input from [1].
Importantly, the “aligned” input direction that we used to demonstrate low-rank suppression in Figure 6a is defined in terms of the dominant right singular vector of the effective connectivity matrix, $W$, i.e., the Jacobian at the fixed point. Using the left singular vector of the explicit connectivity, $Q$, in place of the effective connectivity, $W$, does not produce suppression (see Figure 19c,d in Appendix F). This observation validates our linearization argument and highlights the necessity of using effective connectivity in place of explicit connectivity in nonlinear networks.
@@
Taken together, our results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks can be more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure. For example, if the forcing term is taken to represent interventions like masking or inoculation, then our results predict that it is more effective to apply these interventions randomly to individuals than to apply interventions to individual sub-populations or other structures reflected in the dominant structure of the proximity matrix.
VIII Discussion
@@
We presented theory and examples showing that strongly low-rank recurrent networks close to a stable equilibrium suppress inputs or perturbations aligned with the dominant directions of their connectivity matrices. We also derived conditions under which these networks generate high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
@@
Our results also rely on a linearization of the dynamics around a stable fixed point. The dimensionality of nonlinear network dynamics away from a stable equilibrium is also an important question. However, nonlinear networks can produce any dynamics in response to any inputs [37, 38] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks. Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then used bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations. Combining our approaches by incorporating external inputs into their analytical approach is a promising direction for future work.
We focused on the response of low-rank networks to external input, whereas much previous work focuses on the intrinsic dynamics of low-rank networks in the absence of external drive. Landau and Sompolinsky [4] consider a similar model to ours, but look at the interaction between externally driven responses and intrinsic, chaotic dynamics. However, they restrict to low-dimensional external inputs in the column space of the low-rank part of the connectivity matrix. Future work could combine these approaches to better understand the interaction between chaotic, intrinsic dynamics and high-dimensional external input.
It is tempting to ascribe low-rank suppression to the presence of an eigenvalue with large, negative real part (as in Figure 1 and Figure 3b–d). A large, negative eigenvalue is well-known to produce a suppressive dynamic. However, the results in Figure 2 and Section IV.5 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression. In other words, in recurrent networks driven by external input near a stable equilibrium, low-rank suppression depends only on the presence of large singular values of the connectivity matrix, $W$. High-dimensional dynamics, on the other hand, depend additionally on the singular values of the recurrent alignment matrix, $P=V^{T}U$.
@@
IX Methods
All simulations and analyses were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equations, except for Figure 6 and Figure 19 in Appendix F, were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 6 and Figure 19 in Appendix F were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Networks in Figures 1 and 2 used the linear model from Eq. (1). Networks in Figures 35 used the nonlinear model from Eq. (7). Code to produce all figures is available at